# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v2

@@ 1 Introduction
On the other hand, current DNNs, including the PSENNs discussed above, are generally optimized with the aim of finding the point estimate values of the parameters that minimize a given loss term, leading to models with a deterministic inference process. An important drawback of these approaches is that, generally, it is not possible to measure to what extent the model is confident in its own prediction, known in the literature as the epistemic uncertainty of the model [19, 24, 12]. Indeed, it has been shown that DNNs tend to be overconfident in their predictions [16], classifying inputs with high probability even when random or out-of-distribution inputs are provided. Being unable to detect when the model is making such unreliable or uninformed decisions dramatically reduces the trustworthiness and reliability of the model, and is, therefore, of great concern in high-stakes applications, as recognized in the EU AI Act [11]. To address this issue, recent works proposed replacing point estimates for the parameters of the model with a probability distribution over their values [15]. This leads to a stochastic model, for which it is possible to measure the aforementioned types of uncertainty, and, thus, identify when the model is making uncertain predictions that should not be trusted.
@@
    The effectiveness of Prob-PSENNs is evaluated in several classification tasks, demonstrating their reliability and trustworthiness.
@@ 3 Probabilistic Prototype-Based Self-Explainable Neural Networks (Prob-PSENNs)
As mentioned above, PSENNs treat the prototypes as network parameters, for which point estimate values are computed and hence are fixed once the training is completed. In this section, we introduce Prob-PSENNs, which generalize the prototypes by means of defining and learning a probability distribution over their values, as we illustrate in Figure 1. The main motivations for this probabilistic formulation are discussed below.
First, for each class of the problem, there might exist several sets of prototypes $R$ capable of representing the data with comparable effectiveness. For illustration, let us consider a handwritten digit classification scenario, and the fact that there are several ways of representing each digit. This implies that there are several representative prototypes for each class. Consequently, instead of relying on a single set of prototypes, it might be more reasonable to consider all those possible variations in order to classify an input. Furthermore, in previous approaches, the number of prototypes $m$ needs to be finite and defined beforehand, even if, in practice, it might be cumbersome or infeasible to determine how many prototypes will be required to capture sufficiently diverse characterizations of the classes. In contrast, our probabilistic reformulation enables a more flexible way to decide the variety of the prototypes required to represent the classes, which will be determined, assuming one prototype distribution per class, by the variance of the distributions.
Finally, the choice of the prototypes in $R$ conditions not only the output but also the explanation. Therefore, placing a probability distribution over the prototypes allows us to enable a novel probabilistic framework for the generation of the explanation as well, which can be employed to capture different sources of uncertainty in the explanations of the model, and thereby, produce more informative and useful explanations.
@@ 3.2 Capturing the predictive uncertainty
The predictive distribution defined in Equation (1) allows us to better model the predictive uncertainty of a Prob-PSENN compared to its deterministic counterpart. First, the total predictive uncertainty can be measured by means of the entropy [36] of the prediction: ${H[p(\bm{y}|x)]=-\sum_{i=1}^{c}p(\bm{y}=y_{i}|x)\cdot\log p(\bm{y}=y_{i}|x)}$· However, for a model with fixed (pointwise) values for the parameters, the source of this uncertainty will be only aleatoric uncertainty, which refers to the possibly unavoidable noise, randomness or ambiguity inherent in the data. Considering a probability distribution for the parameter values instead of point estimates allows us to also capture the epistemic uncertainty [19, 21, 24], which accounts for the model uncertainty about the true parameters modeling the data (e.g., as a consequence of insufficient training data). This enables a safer use of the model in critical tasks, since it allows us to identify when the model is making unreliable predictions. An illustrative comparison between aleatoric and epistemic uncertainty is provided in Figure 2 (left).
@@ 3.3.2 Quantifying explanation uncertainty
Regarding epistemic explanatory uncertainty, as explained above, high uncertainty will be achieved for those inputs for which the distances to all the prototypes are large. On that account, the minimum sample mean: $\bar{\delta}_{x,i*}=\min_{1\leq j\leq c}\bar{\delta}_{x,j}$ represents a suitable base metric for evaluating this uncertainty. Furthermore, in order to determine the severity of the uncertainty in a normalized and scale-independent manner, a quantile-based metric will be defined, close to the approach proposed in [33]. This metric will be based on comparing the value at hand with the values obtained for the data distribution $\mathscr{D}$, which can be estimated, in practice, using a training set $D$:
@@ 3.5 Training procedure
Let $p_{\lambda}(\bm{R})$ represent a parametric probability distribution over the prototypes. The goal of the training procedure will be to optimize the parameters $\lambda$, jointly with the rest of parameters of the model, in order to maximize both the predictive performance and interpretability of the model. For simplicity, the prototypes $\bm{r}_{1}\dots,\bm{r}_{c}$ will be assumed to be independent random variables, with each $\bm{r}_{i}$ distributed according to a distribution $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$, and, hence, $p_{\lambda}(\bm{R})=p_{\lambda}(\bm{r}_{1},\dots,\bm{r}_{c})=\prod_{i=1}^{c}p_{\lambda_{i}}(\bm{r}_{i})$. The following generalized loss function will be employed to train the entire architecture:
@@ 4.1 Setup
The effectiveness and capabilities of the proposed model will be assessed on the MNIST [26], Fashion-MNIST [44], K-MNIST [8] and SVHN [31] datasets. This choice of datasets is based on the fact that they provide a clear and coherent notion of prototypes to facilitate the assessment of our results, particularly regarding explanation quality (e.g., prototype diversity) and uncertainty quantification (e.g., ambiguity).
The autoencoder module of Prob-PSENN will be implemented by a CNN-based Variational AutoEncoder (VAE), which adds an additional regularization term to the overall loss function, namely, the Kullback–Leibler (KL) divergence between the learned latent distribution and a prior (Gaussian) distribution [23]. The full implementation details are reported in Appendix B. For each prototype distribution $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$, a full covariance Gaussian distribution will be assumed, and, as in [27], the squared Euclidean distance will be used as the distance metric $\phi$ to measure the distance between the encoded input and the sampled prototypes (in the latent space). The final classification layer $W$ will be set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal, further promoting transparency. The Adam optimization method [22] has been used to train the networks, using a learning rate of $0.001$. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
@@ 4.4 Comparison between conventional PSENN and Prob-PSENN
To analyze the predictive performance of Prob-PSENNs, we begin by comparing the accuracies obtained for the MNIST, F-MNIST and K-MNIST test sets, as shown in Table 1. For reference, we include the results of the deterministic baseline models proposed in Li et al. [27] and Gautam et al. [13], reproduced using the implementation released by the authors (see Appendix A for a further discussion on these works). The results show that Prob-PSENNs achieve either superior or competitive performance in all the datasets, while having as benefits greater robustness, trustworthiness, and enhanced explanatory capabilities. These capabilities are showcased in Figure 10, which compares, for the MNIST task, the explanation computed by a conventional PSENN ($m=15$) with that computed by Prob-PSENN. More comparisons are provided in Appendices D and E. As can be seen, the prototypes sampled by Prob-PSENN exhibit considerably more diversity than the fixed prototypes learned by a conventional PSENN, which enables one to find more tailored prototypes for the input at hand, as well as to provide alternative forms that the class might take. In addition, while PSENNs only return point distances to the prototypes, Prob-PSENNs provide a distribution over those distances, enabling assessment of the uncertainty of the prediction, and, hence, accounting for a more trustworthy approach.
@@ 4.4.2 Ablation studies
In addition, Appendix F further examines the benefits of using a VAE over a vanilla AE as the backbone architecture. Table F.2 compares the predictive performance of Prob-PSENNs when replacing the VAE with an AE while keeping all other architectural and training details unchanged. The results show that AEs lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.21 visualizes randomly sampled prototypes from Prob-PSENNs with different backbone architectures, showing that while vanilla AE backbones provide comparable results in lower latent dimensions (e.g., $l=10$), they struggle to maintain high-quality prototype distributions as the dimensionality increases. In contrast, VAEs enable effective scaling to higher-dimensional spaces. This demonstrates that, while Prob-PSENNs are not inherently restricted to VAE backbones, more regularized latent spaces enhance both predictive performance and prototype quality in high-dimensional settings.
@@ 4.5 Scalability
In order to show that our approach can scale to larger numbers of classes, Appendix G includes results on the E-MNIST dataset [9], an extension of MNIST which contains handwritten representations of numbers and letters and is composed of 47 classes. As can be seen, Prob-PSENN achieves a highly diverse and representative sets of prototypes for all the classes, which would require an excessively large number of pointwise prototypes to achieve using deterministic approaches. Furthermore, to show that Prob-PSENNs can scale to colour datasets, we tested our model on the SVHN dataset. The accuracy of the model in this dataset is reported in Table 3. The results show that the predictive performance of Prob-PSENNs is on par with previous deterministic approaches, when not superior.
@@ 5 Conclusions
In this paper we introduced Prob-PSENN, a novel Prototype-Based Self-Explainable Neural Network paradigm relying on a probabilistic generalization of the prototypes, which enhances the capabilities of the model in two key aspects. First, unlike previous approaches, our model is not restricted to a fixed number or set of prototypes. This enables Prob-PSENN to learn much more diverse prototypes, and to consider all of them when classifying an input, which represents a more flexible and robust approach. Second, our probabilistic approach equips Prob-PSENNs with tools to capture the uncertainty in both the prediction and the explanation of the network. This novel feature makes it possible to detect when the model is providing uncertain outcomes, substantially increasing the reliability of the model. We experimentally demonstrated how Prob-PSENNs provide more informative and trustworthy explanations than their non-probabilistic counterparts, while achieving superior or competitive predictive accuracies. Moreover, we show that the accuracy of Prob-PSENNs can be improved by leveraging their novel uncertainty-quantification capabilities, further enhancing their reliability.
@@ 6 Limitations and future work
As future work, inspired by recent advances in Bayesian Neural Networks [30, 12, 15], we plan to develop a Bayesian formulation for the distribution over the prototypes, which might enable a more theoretically-grounded approach for uncertainty quantification. Furthermore, we also believe that a fully-probabilistic model (i.e., placing a probability distribution on the rest of the network parameters as well) could enhance the quantification of the uncertainty in Prob-PSENNs, or even induce new notions of uncertainty, such as the encoding or class-representation uncertainty. Apart from that, while Prob-PSENNs have demonstrated high predictive and explanatory performance, we found it challenging to model probability distributions over the prototypes in very high-dimensional spaces, and, therefore, resort to relatively low dimensional latent spaces in our experiments. We plan to address this limitation in the future, in order to scale the proposed approach to higher-dimensional datasets.

### v3

@@ 1 Introduction
On the other hand, current DNNs, including the PSENNs discussed above, are generally optimized with the aim of finding the point estimate values of the parameters that minimize a given loss term, leading to models with a deterministic inference process. An important drawback of these approaches is that, generally, it is not possible to measure to what extent the model is confident in its own prediction, known in the literature as the epistemic uncertainty of the model [20, 25, 13]. Indeed, it has been shown that DNNs tend to be overconfident in their predictions [17], classifying inputs with high probability even when random or out-of-distribution inputs are provided. Being unable to detect when the model is making such unreliable or uninformed decisions dramatically reduces the trustworthiness and reliability of the model, and is, therefore, of great concern in high-stakes applications, as recognized in the EU AI Act [12]. To address this issue, recent works proposed replacing point estimates for the parameters of the model with a probability distribution over their values [16]. This leads to a stochastic model, for which it is possible to measure the epistemic uncertainty, and, thus, identify when the model is making uncertain predictions that should not be trusted.
@@
    The effectiveness of Prob-PSENNs is evaluated on several image and tabular data classification tasks, demonstrating their reliability and trustworthiness.
@@ 3 Probabilistic Prototype-Based Self-Explainable Neural Networks (Prob-PSENNs)
As mentioned above, PSENNs treat the prototypes as network parameters, for which point estimate values are computed and hence are fixed once the training is completed. In this section, we introduce Prob-PSENNs, which generalize the prototypes by means of defining and learning a probability distribution over their values, as we illustrate in Figure 1. The main motivation for this probabilistic formulation is discussed below.
First, for each class of the problem, there might exist several sets of prototypes $R$ capable of representing the data with comparable effectiveness. For illustration, let us consider a handwritten digit classification scenario, and the fact that there are several ways of representing each digit. This implies that there are several representative prototypes for each class. Consequently, instead of relying on a single set of prototypes, it might be more reasonable to consider all those possible variations in order to classify an input. Furthermore, in previous approaches, the number of prototypes $m$ needed to be finite and defined beforehand, even if, in practice, it might be cumbersome or infeasible to determine how many prototypes will be required to capture sufficiently diverse characterizations of the classes. In contrast, our probabilistic reformulation enables a more flexible way to decide the variety of the prototypes required to represent the classes, which will be determined, assuming one prototype distribution per class, by the variance of the distributions.
Finally, the choice of the prototypes in $R$ conditions not only the output but also the explanation. Therefore, placing a probability distribution over the prototypes allows us to enable a novel probabilistic framework for the generation of the explanation as well, which can be employed to capture different sources of uncertainty in the explanations of the model, and thereby produce more informative and useful explanations.
@@ 3.2 Capturing the predictive uncertainty
The predictive distribution defined in Equation (1) allows us to better model the predictive uncertainty of a Prob-PSENN compared to its deterministic counterpart. First, the total predictive uncertainty can be measured by means of the entropy [37] of the prediction: ${H[p(\bm{y}|x)]=-\sum_{i=1}^{c}p(\bm{y}=y_{i}|x)\cdot\log p(\bm{y}=y_{i}|x)}$. However, for a model with fixed (pointwise) values for the parameters, the source of this uncertainty will be only aleatoric uncertainty, which refers to the possibly unavoidable noise, randomness or ambiguity inherent in the data. Considering a probability distribution for the parameter values instead of point estimates allows us to also capture the epistemic uncertainty [20, 22, 25], which accounts for the model uncertainty about the true parameters modeling the data (e.g., as a consequence of insufficient training data). This enables a safer use of the model in critical tasks, since it allows us to identify when the model is making unreliable predictions. An illustrative comparison between aleatoric and epistemic uncertainty is provided in Figure 2 (left).
@@ 3.3.2 Quantifying explanation uncertainty
Regarding epistemic explanatory uncertainty, as explained above, high uncertainty will be achieved for those inputs for which the distances to all the prototypes are large. On that account, the minimum sample mean: $\bar{\delta}_{x,i*}=\min_{1\leq j\leq c}\bar{\delta}_{x,j}$ represents a suitable base metric to evaluate this uncertainty. Furthermore, in order to determine the severity of the uncertainty in a normalized and scale-independent manner, a quantile-based metric will be defined, close to the approach proposed in [34]. This metric will be based on comparing the value at hand with the values obtained for the data distribution $\mathscr{D}$, which can be estimated, in practice, using a training set $D$:
@@ 3.5 Training procedure
Let $p_{\lambda}(\bm{R})$ represent a parametric probability distribution over the prototypes. The goal of the training procedure will be to optimize the parameters $\lambda$, jointly with the rest of the parameters of the model, in order to maximize both the predictive performance and interpretability of the model. For simplicity, the prototypes $\bm{r}_{1}\dots,\bm{r}_{c}$ will be assumed to be independent random variables, with each $\bm{r}_{i}$ distributed according to a distribution $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$, and, hence, $p_{\lambda}(\bm{R})=p_{\lambda}(\bm{r}_{1},\dots,\bm{r}_{c})=\prod_{i=1}^{c}p_{\lambda_{i}}(\bm{r}_{i})$. The following generalized loss function will be employed to train the entire architecture:
@@ 4.1 Setup
The effectiveness and capabilities of the proposed model will be assessed mainly on the MNIST [27], Fashion-MNIST [45], Kuzushiji-MNIST (also known as K-MNIST) [9] and SVHN [32] datasets. This choice of datasets is based on the fact that they provide a clear and coherent notion of prototypes to facilitate the assessment of our results, particularly regarding explanation quality (e.g., prototype diversity) and uncertainty quantification (e.g., ambiguity), while offering increasing complexity in the representations (from digits to visually complex handwritten Hiragana characters). Additional experiments will be carried out for the EMNIST [10] dataset, in order to assess the scalability of our model to a larger number of classes, and the Sensorless Drive Diagnosis dataset [6], in order to illustrate the applicability of Prob-PSENN to tabular data.
The autoencoder module of Prob-PSENN will be implemented by a CNN-based Variational AutoEncoder (VAE) for image datasets, which adds an additional regularization term to the overall loss function, namely, the Kullback–Leibler (KL) divergence between the learned latent distribution and a prior (Gaussian) distribution [24]. For tabular data, an MLP architecture will be employed for the autoencoder. The full implementation details are reported in Appendix B. For each prototype distribution $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$, a full covariance Gaussian distribution will be assumed, whose parameters (mean vectors and covariance matrices) are optimized jointly with the rest of the network parameters. Covariance matrices are parameterized via Cholesky decomposition to ensure positive definiteness. As in [28], the squared Euclidean distance will be used as the distance metric $\phi$ to measure the distance between the encoded input and the sampled prototypes (in the latent space). The final classification layer $W$ will be set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal, further promoting transparency. The Adam optimization method [23] has been used to train the networks, using a learning rate of $0.001$. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
@@ 4.4 Comparison between conventional PSENN and Prob-PSENN
To analyze the predictive performance of Prob-PSENNs, we begin by comparing the accuracies obtained for the MNIST, F-MNIST and K-MNIST test sets, as shown in Table 1. For reference, we include the results of the deterministic baseline models proposed in Li et al. [28] and Gautam et al. [14], reproduced using the implementation released by the authors (see Appendix A for a further discussion on these works). Results on the SVHN dataset will be reported in Table 3 (see Section 4.5), using a complementary set of baselines more competitive for color images. The results show that Prob-PSENNs achieve either superior or competitive performance in all the datasets, while having as benefits greater trustworthiness and enhanced explanatory capabilities. These capabilities are showcased in Figure 10, which compares, for the MNIST task, the explanation computed by a conventional PSENN ($m=15$) with that computed by Prob-PSENN. More comparisons are provided in Appendices D and E. As can be seen, the prototypes sampled by Prob-PSENN exhibit considerably more diversity than the fixed prototypes learned by a conventional PSENN, which enables one to find more tailored prototypes for the input at hand, as well as to provide alternative forms that the class might take. In addition, while PSENNs only return point distances to the prototypes, Prob-PSENNs provide a distribution over those distances, enabling assessment of the uncertainty of the prediction, and, hence, accounting for a more trustworthy approach.
@@ 4.4.2 Ablation studies
Furthermore, Table F.1 in Appendix F also compares Prob-PSENN with black-box baselines that share the same encoder architecture, followed by a linear classifier trained end-to-end with cross-entropy loss. To ensure a fair comparison, we also evaluated a variant using the full VAE backbone (encoder and decoder), matching the regularization and training dynamics of Prob-PSENN’s autoencoder, even if the decoder plays no functional role in the black-box classification process at inference time. Consistent with prior work [28, 14], our results show that Prob-PSENN matches or outperforms these black-box counterparts. We attribute this to the structured clustering induced by the prototype distributions in the latent space and the robustness gained through averaging predictions over multiple sampled prototypes, reinforcing the practical advantages of Prob-PSENN’s pipeline.
In addition, Appendix F further examines the benefits of using a VAE over a vanilla AE as the backbone architecture. Table F.2 in the appendix compares the predictive performance of Prob-PSENNs when replacing the VAE with an AE while keeping all other architectural and training details unchanged. The results show that AEs lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.22 (Appendix F) visualizes randomly sampled prototypes from Prob-PSENNs with different backbone architectures, showing that while vanilla AE backbones provide comparable results in lower latent dimensions (e.g., $l=10$), they struggle to maintain high-quality prototype distributions as the dimensionality increases. In contrast, VAEs enable effective scaling to higher-dimensional spaces. This demonstrates that, while Prob-PSENNs are not inherently restricted to VAE backbones, more regularized latent spaces enhance both predictive performance and prototype quality in high-dimensional settings.
@@ 4.5 Scalability
In order to show that our approach can scale to larger numbers of classes, Appendix G includes results on the E-MNIST dataset [10], an extension of MNIST which contains handwritten representations of numbers and letters and is composed of 47 classes. As can be seen, Prob-PSENN achieves highly diverse and representative sets of prototypes for all the classes, which would require an excessively large number of pointwise prototypes to achieve using deterministic approaches. Furthermore, to show that Prob-PSENNs can scale to colour datasets, we tested our model on the SVHN dataset. The accuracy of the model in this dataset is reported in Table 3. The results show that the predictive performance of Prob-PSENNs is on par with previous deterministic approaches, when not superior.
4.6 Application to tabular data
Although prototype-based self-explainable neural networks are often evaluated only on image data [28, 14, 8, 18, 21, 44], in this section, we show that Prob-PSENN can be applied naturally to tabular settings. We demonstrate this using the Sensorless Drive Diagnosis (SDD) dataset [6], which contains 58,509 examples of electric motor signals represented by 48 real-valued features and classified into 11 operating conditions. We trained Prob-PSENN using the same pipeline described for images, leveraging a three-layer MLP as an encoder with batch normalization and LeakyReLU activations. The decoder mirrors this structure in reverse.
Figure 12 shows an example explanation produced by a Prob-PSENN trained on the SDD dataset. Given the non-visual nature of the data, we leverage a domain-appropriate visualization strategy based on radar plots to visualize the input and the prototypes. The top row includes the radar plot corresponding to the input sample, alongside the predictive probability distribution over the classes. The three remaining rows include prototypes corresponding to the three most likely classes, in descending order of confidence. Each of these rows contains the three closest prototypes on the left and the farthest prototype on the right. Notice that each prototype is visualized as a color radar chart (solid lines), with the shape of the input overlaid for comparison (dashed lines). The overlapping areas reveal which sensor patterns in the original input match the class fingerprints (large overlap) and which ones deviate (visible gaps).
As can be seen, the explanation highlights both prototypical and divergent feature patterns, helping to understand the input classification, which patterns are shared with other classes, and which features drive uncertainty or ambiguity. These visual explanations show how Prob-PSENN can capture class-representative structures even in tabular domains, broadening the applicability of our model.
Figure 12: Explanation of a prediction made by Prob-PSENN on an input from the Sensorless Drive Diagnosis tabular dataset. The first row contains the input (left) and the class probabilities predicted by Prob-PSENN (middle). The next three rows represent class prototypes sampled from Prob-PSENN, considering, for simplicity, the three most likely classes. Each of these rows displays the three closest prototypes (left) and the most distant one (right), visualized as colored radar plots. The outline of the input (black, dashed line) is overlaid on each prototype radar plot, for direct comparison. Overlaps between the input and class prototypes highlight shared patterns, while mismatches reveal deviations or missing features, offering insights into the model’s reasoning, uncertainty, and potential counterfactuals.
@@ 5 Conclusions
In this paper we introduced Prob-PSENN, a novel Prototype-Based Self-Explainable Neural Network paradigm relying on a probabilistic generalization of the prototypes, which enhances the capabilities of the model in two key aspects. First, unlike previous approaches, our model is not restricted to a fixed number or set of prototypes. This enables Prob-PSENN to learn much more diverse prototypes, and to consider all of them when classifying an input, which represents a more flexible and robust approach. Second, our probabilistic approach equips Prob-PSENNs with tools to capture the uncertainty in both the prediction and the explanation of the network. This novel feature makes it possible to detect when the model is providing uncertain outcomes, substantially increasing the reliability of the model. We experimentally demonstrated how Prob-PSENNs provide more informative and trustworthy explanations than their non-probabilistic counterparts, while achieving superior or competitive predictive accuracies. Moreover, we show that the accuracy of Prob-PSENNs can be improved by leveraging their novel uncertainty-quantification capabilities, further enhancing their reliability. Finally, we demonstrate that Prob-PSENN extends naturally to tabular data, enabling interpretable decision-making beyond vision tasks.
@@ 6 Limitations and future work
As future work, inspired by recent advances in Bayesian Neural Networks [31, 13, 16], we plan to develop a Bayesian formulation for the distribution over the prototypes, which might enable a more theoretically-grounded approach for uncertainty quantification. Furthermore, we also believe that a fully-probabilistic model (i.e., placing a probability distribution on the rest of the network parameters as well) could enhance the quantification of the uncertainty in Prob-PSENNs, or even induce new notions of uncertainty, such as the encoding or class-representation uncertainty. Apart from that, while Prob-PSENNs have demonstrated high predictive and explanatory performance, we found it challenging to model probability distributions over the prototypes in very high-dimensional spaces, and, therefore, resort to relatively low dimensional latent spaces in our experiments. We plan to address this limitation in the future, in order to scale the proposed approach to higher-dimensional datasets. Finally, in future work, we plan to extend Prob-PSENN to more complex data modalities, including discrete or heterogeneous tabular datasets that combine categorical and numerical features, as well as explore its applicability to other structured and unstructured data domains.